# Basis-Spline Assisted Coded Computing: Strategies and Error Bounds

### v1

@@ Abstract
Coded computing has become a key framework for reliable distributed computation over decentralized networks, effectively mitigating the impact of stragglers. Although there exists a wide range of coded computing methods to handle both polynomial and non-polynomial functions, computing methods for the latter class have received traction due its inherent challenges in reconstructing non-polynomial functions using a finite number of evaluations. Among them, the state-of-the-art method is Berrut Approximated coded computing, wherein Berrut interpolants, are used for approximating the non-polynomial function. However, since Berrut interpolants have global support characteristics, such methods are known to offer degraded accuracy when the number of stragglers is large. To address this challenge, we propose a coded computing framework based on cubic B-spline interpolation. In our approach, server-side function evaluations are reconstructed at the master node using B-splines, exploiting their local support and smoothness properties to enhance stability and accuracy. We provide a systematic methodology for integrating B-spline interpolation into coded computing and derive theoretical bounds on approximation error in terms of the number of servers and stragglers. Comparative analysis demonstrates that our framework significantly outperforms Berrut-based methods for various non-polynomial functions.
@@ I Introduction
Distributed computing is an effective paradigm for executing large-scale computational tasks across multiple parallel computing machines. In such systems, the data sets is distributed among several worker nodes, while a master node coordinates the computation and aggregates the results [1]. A major performance bottleneck in distributed learning systems is the presence of stragglers, i.e., slow or unresponsive worker nodes, which can significantly delay the overall computation. To address this challenge, coded computing has been proposed as a fault-tolerant and efficient approach that mitigates the impact of stragglers by introducing structured redundancy into distributed computations. In particular, by distributing encoded shares of the data among worker nodes, coded computing enables the master node to recover the desired result from the outputs of only a subset of workers, thereby avoiding the need to wait for stragglers. This idea has been successfully applied to problems such as distributed matrix multiplication [2, 3] and polynomial function computation [4, 5, 6]. However, many existing coded computing schemes such as Lagrange Coded Computing(LCC) [4], Analog Lagrange Coded Computing(ALCC) [6] are limited to polynomial computations and rely on exact interpolation, which can be numerically unstable over real-valued data [7] and often require a large number of worker nodes. In contrast, [8] proposed a numerically stable coded computing framework based on Chebyshev polynomials; however, this approach remains restricted to polynomial computations and due to its reliance on exact interpolation, still requires a large number of workers as the degree of polynomial increases.
To address these challenges, Berrut Approximated Coded Computing (BACC) [9] was recently proposed as a numerically stable framework for approximately computing arbitrary (possibly non-polynomial) functions in a distributed setting. BACC employs Berrut rational function for both data encoding and decoding, thereby extending existing coded computing frameworks beyond polynomial functions while providing provable approximation error guarantees. Although BACC substantially improves numerical stability, its decoding procedure relies on global approximation, which limits the achievable approximation order as the number of stragglers increases. These limitations motivate the need to investigate new coded computing designs that improve numerical accuracy while explicitly characterizing the associated trade-offs with respect to computation, communication, and straggler tolerance. In parallel, Basis splines have been extensively studied in approximation theory and numerical analysis as a powerful tool for interpolation and function approximation [10, 11][12][13, 14]. Since the foundational work of Schoenberg, B-splines have been shown to possess several desirable properties, including local support, high-order approximation capability, and strong numerical stability. Despite this rich theoretical foundation, spline-based methods have not been systematically explored in the context of coded computing, where existing approaches continue to rely predominantly on polynomial or rational interpolation schemes.
@@
In this work, we bridge this gap by proposing a B-spline assisted coded computing (BSCC) framework for distributed computation in the presence of stragglers. Unlike global interpolation-based schemes, BSCC employs B-spline basis functions to perform local, piecewise-polynomial approximation of the target function, enabling higher-order approximation accuracy and improved numerical stability. Our specific contributions are listed below:
@@
1) For coded computing frameworks, we formulate a reconstruction methodology based on clamped B-spline interpolation, wherein the composite function evaluations returned by the workers are represented using B-spline basis functions. Our methodology explicitly specifies the construction of the spline space, the use of clamped knot vectors, and the incorporation of natural boundary conditions, resulting in a well-posed linear system for reconstruction at the master node. We analyze the structure of the resulting B-spline interpolation system, and demonstrate that the reconstruction can be carried out efficiently using banded linear solvers, with computational complexity that scales linearly with the number of workers.
2) We establish explicit bounds on the Lebesgue constant associated with the proposed spline-based decoding scheme and connect these bounds to the overall approximation error of the distributed computation. This analysis yields a precise characterization of how the reconstruction error depends on the number of straggling workers.
3) Through both theoretical comparison and numerical simulations, we show that the proposed B-spline–based coded computing framework achieves strictly better approximation accuracy than Berrut Approximated Coded Computing (BACC), exhibiting significantly faster error decay as the number of stragglers increases (see Section IV).
@@ II Preliminaries on B-Splines
In this subsection, we revisit some basic definitions and fundamental properties of B-spline functions.
Definition 1 (Clamped (open) knot vector[11]).
Let $a,b]\subset\mathbb{R}$ be a compact interval, and let $\{\alpha_{i}\}_{i=0}^{N-1}\subset[a,b]$ be a strictly increasing sequence of evaluation points, i.e., $a=\alpha_{0}<\alpha_{1}<\cdots<\alpha_{N-1}=b$. For a given spline degree $p\geq 0$, spline interpolation of $N$ data points requires a non-decreasing knot sequence of length $N+2p$. A knot sequence is said to be $(p+1)$-open (or clamped) on $[a,b]$ if it is non-decreasing and has end knots of multiplicity $p+1$, i.e., $a=t_{0}=\cdots=t_{p}<t_{p+1}\leq\cdots\leq t_{N+p-2}<t_{N+p-1}=\cdots=t_{N+2p-1}=b$. Throughout this work, we employ the $(p+1)$-open (clamped) knot vector $t=(t_{j})_{j=0}^{N+2p-1}$ defined by
$t=[\underbrace{\alpha_{0},\ldots,\alpha_{0}}_{p+1},\alpha_{1},\alpha_{2},\ldots,\alpha_{N-2},\underbrace{\alpha_{N-1},\ldots,\alpha_{N-1}}_{p+1}].$ (1)
Definition 2 (B-spline basis [10]).
Given the knot vector $t$ and spline degree $p\geq 0$, the B-spline basis functions $\{B_{j,p}\}_{j=0}^{N+p-2}$ associated with $\mathcal{S}_{p,t}$ are defined recursively by the Cox–de Boor formula. For $p=0$,
@@
Definition 3 (Support and smoothness of B-splines[12]).
Each B-spline basis function $B_{j,p}$ belongs to $C^{p-1}(a,b)$ and has compact support $\operatorname{supp}(B_{j,p})=(t_{j},t_{j+p+1})$. Moreover, $B_{j,p}(z)\geq 0$ for all $z\in(t_{j},t_{j+p+1})$. Consequently, each basis function is nonzero only over $p+1$ consecutive knot intervals.
Definition 4 (Spline space).
Let $t$ be the clamped knot vector defined in Definition 1. The spline space $\mathcal{S}_{p,t}$ consists of all piecewise polynomial functions of degree at most $p$ defined on $[a,b]$ with smoothness determined by the multiplicities of the knots.
Remark 1 (Effect of knot multiplicity[11]).
For a spline of degree $p$, a knot of multiplicity $m$ induces $C^{p-m}$ smoothness at that knot. In particular, the boundary knots in the clamped knot vector have multiplicity $p+1$, resulting in $C^{-1}$ regularity at the endpoints, while interior knots of unit multiplicity ensure $C^{p-1}$ smoothness.
Definition 5 (B-spline representation of spline interpolant [12]).
Let $f\in C[a,b]$ be a given function, and let $\{\alpha_{i}\}_{i=0}^{N-1}\subset[a,b]$ be a strictly increasing set of evaluation points. Let $t$ be a prescribed knot vector and let $\mathcal{S}_{p,t}:=\operatorname{span}\{B_{0,p},B_{1,p},\ldots,B_{N+p-2,p}\}$ denote the associated spline space of degree $p$, where $\{B_{j,p}\}_{j=0}^{N+p-2}$ are the corresponding B-spline basis functions. A spline interpolant $s\in\mathcal{S}_{p,t}$ is expressed as
$s(z)=\sum_{j=0}^{N+p-2}C_{j,p}\,B_{j,p}(z),$ (2)
where the coefficients $\{C_{j,p}\}$ are determined by enforcing the interpolation conditions $s(\alpha_{i})=f(\alpha_{i}),\quad i=0,\ldots,N-1.$ Substituting (2) into the interpolation conditions yields the linear system
$\sum_{j=0}^{N+p-2}C_{j,p}\,B_{j,p}(\alpha_{i})=f(\alpha_{i}),\qquad i=0,\ldots,N-1.$ (3)
@@ III B-Spline Assisted Coded Computing
We propose a B-spline assisted coded computing (BSCC) framework for approximate computation of an arbitrary function $f(\cdot)$ over a given input dataset. Our setup consists of a master node and $N$ worker nodes indexed by the set $\mathcal{W}=\{\mathcal{W}_{0},\mathcal{W}_{1},\ldots,\mathcal{W}_{N-1}\}$. We assume that each worker node communicates exclusively with the master node through a dedicated communication link, and no communication is permitted among workers. The input dataset is stored at the master node, while all workers are assumed to have prior knowledge of the function to be evaluated. Among the worker nodes, we assume the presence of $s$ number of straggling workers, for $1\leq s<N-2$, which may fail to return their computation results within the prescribed time. Let $\mathbf{X}=(\mathbf{X}_{0},\ldots,\mathbf{X}_{K-1})$ denote the dataset held by the master node, where $\mathbf{X}_{j}\in\mathbb{R}^{m\times n}$ for all $j\in[K]$, with $K]\triangleq\{0,1,2,\ldots,K-1\}$. The primary objective of BSCC framework is to approximately evaluate an arbitrary function $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ on the matrices in $\mathbf{X}$ in a decentralized manner while guaranteeing tolerable approximation errors against $s$ number of stragglers. In particular, since the function $f(\cdot)$ may be non-polynomial, BSCC aims to enable numerically stable distributed computation of $f(\cdot)$ on $\mathbf{X}_{j}$ for all $j\in[K]$ with bounded approximation error; that is, if $\mathbf{Y}_{j}$ denotes the output reconstructed from the distributed computation corresponding to $\mathbf{X}_{j}$, then $\mathbf{Y}_{j}\approx f(\mathbf{X}_{j})$ for all $j\in[K]$ within a tolerable accuracy loss. Within this setting, we employ a generalized encoding method available in the literature and propose a B-spline assisted numerically stable reconstruction procedure at the master. [^1] Coded computing under additional constraints such as data privacy and Byzantine adversaries are well studied. In this work, we restrict our attention to resilience against straggling workers. Privacy guarantees and Byzantine robustness are straightforward to generalize using existing techniques.
@@ III-A Encoding
We describe the encoding method, specifically highlighting how the dataset $\mathcal{X}$ is mapped into distributed shares for the worker nodes. To encode the dataset $\mathcal{X}$, the master node may employ any suitable interpolatory basis functions $\{\Phi_{j}(\cdot)\}$, such as the Lagrange basis [6] or the Berrut basis functions [9]. These basis functions satisfy the cardinality property, namely $\Phi_{j}(\beta_{i})=1$ for $j=i$ and $\Phi_{j}(\beta_{i})=0$ for $j\neq i$, for the underlying $K$ ordered non-decreasing encoding points $\{\beta_{i}\in\mathbb{R}~|~i=0,1,\ldots,K-1\}$. Using this property, the dataset $\mathcal{X}$ is encoded as a function of an evaluation variable $z$ as
@@
III-B Distribution of Shares among the Worker Nodes
Once $u(z)$ is constructed as in (4), the master selects a set of $N$ distinct and ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}\subset\mathbb{R}$, such that $\alpha_{0}<\alpha_{1}<\cdots<\alpha_{N-1}$. For instance, these points could be chosen as Chebyshev nodes of the first kind or second kind. Using these evaluation points, the master node generates the encoded shares by evaluating the encoding function $u(z)$ at each $\alpha_{i}$, resulting in the set of function values $\{\mathbf{Y}_{i}\}_{i=0}^{N-1}$, where $\mathbf{Y}_{i}=u(\alpha_{i}),\quad i\in[N-1].$ Finally, the master distributes the encoded shares to the worker nodes by assigning the share $\mathbf{Y}_{i}$ to worker $\mathcal{W}_{i}$, for $i\in[N-1]$.
III-C Computation at the Worker Nodes
Once the worker $\mathcal{W}_{i}$ receives its evaluation $\mathbf{Y}_{i}$, it computes $f(\mathbf{Y}_{i})=f(u(\alpha_{i}))$, and returns it to the master. In contrast, the straggling workers maynot return these values within the prescribed deadline.
III-D Function Reconstruction using Spline
In the absence of stragglers (i.e., when $s=0$), the master node receives $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$ from $N$ worker nodes, corresponding to a set of distinct and ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$. The objective is to reconstruct an approximation of the composite target function $f(u(z))$ over the interval spanned by these points. Since the function $f(\cdot)$ is arbitrary and may be non-polynomial, exact interpolation [6] is generally not feasible or numerically stable. Consequently, we approximate $f(u(z))$ using a basis-spline function $r_{B-spline}(z)\approx f(u(z))$, where

### v2

@@ Abstract
Coded computing has emerged as a key framework for addressing the impact of stragglers in distributed computation. While polynomial functions often admit exact recovery under existing coded computing schemes, non-polynomial functions require approximate reconstruction from a finite number of evaluations, posing significant challenges. Consequently, interpolation-based methods for non-polynomial coded computing have gained attention, with Berrut approximated coded computing emerging as a state-of-the-art approach. However, due to the global support of Berrut interpolants, the reconstruction accuracy degrades significantly as the number of stragglers increases. To address this challenge, we propose a coded computing framework based on cubic B-spline interpolation. In our approach, server-side function evaluations are reconstructed at the master using B-splines, exploiting their local support and smoothness properties to enhance stability and accuracy. We provide a systematic methodology for integrating B-spline interpolation into coded computing and derive theoretical bounds on approximation error for certain class of smooth functions. Our analysis demonstrates that the error bounds of our approach exhibit a faster decay with respect to the number of workers compared to the Berrut-based method. Experimental results also confirm that our method offers improved accuracy over Berrut-based methods for various smooth non-polynomial functions.
@@ I Introduction
Distributed computing is an effective paradigm for executing large-scale computational tasks, wherein a master coordinates the computation by distributing the dataset among several workers [1]. A major performance bottleneck in such systems is the presence of stragglers, i.e., slow or unresponsive workers, which can significantly delay the overall computation. To address this challenge, coded computing has been proposed as a fault-tolerant approach by introducing structured redundancy into distributed computations. This idea has been successfully applied to problems such as distributed matrix multiplication [2, 3] and polynomial function computation [4, 5, 6]. However, many existing coded computing schemes such as Lagrange Coded Computing [4], Analog Lagrange Coded Computing [6] are limited to polynomial computations and rely on exact interpolation, which can be numerically unstable over real-valued data [7] and often require a large number of workers. In contrast, [8] proposed a numerically stable coded computing framework based on Chebyshev polynomials, however, this approach remains restricted to polynomial computations and due to its reliance on exact interpolation, still requires a large number of workers as the degree of polynomial increases.
To address these challenges, Berrut Approximated Coded Computing (BACC) [9] was proposed as a numerically stable framework for approximately computing arbitrary functions. BACC employs Berrut rational functions for both data encoding and reconstruction, thereby extending existing coded computing frameworks beyond polynomial functions while providing provable approximation error guarantees. Although BACC improves numerical stability, its reconstruction procedure relies on global approximation, which limits the achievable approximation order as the number of stragglers increases. To alleviate this problem, [10] explored a learning-theoretic framework to coded computing, wherein the underlying encoding and reconstruction operations were posed as regression problems, and were subsequently solved using data-centric methods. Interestingly, [10] was shown to outperform [9] in terms of relative error when handling non-polynomial functions. Although [10] is the state-of-the art in terms of accuracy metric, it is accompanied by other implementation overheads such as dataset management and model-training. Thus, within the class of non data-centric methods, BACC is still the state-of-the-art, and therefore, there is a need to investigate new structured coded computing methods that outperform numerical accuracy of BACC while explicitly characterizing the associated trade-offs.
I-A Contributions
@@
Basis splines (B-splines) have been extensively studied in approximation theory and numerical analysis as a powerful tool for interpolation and function approximation [11, 12, 13, 14, 15]. In particular, they possess several desirable properties, including local support, higher-order approximation capability, and strong numerical stability. Despite this rich theoretical foundation, spline-based methods have not been systematically explored in the context of coded computing, where existing approaches continue to rely predominantly on polynomial or rational interpolation schemes. To bridge this gap, in this work, we propose a B-spline assisted coded computing (BSCC) framework for distributed computation of arbitrary functions in the presence of stragglers. Unlike global interpolation-based schemes, BSCC employs B-spline basis functions to perform local, piecewise-polynomial approximation of the target function, enabling higher-order approximation accuracy and improved numerical stability.
@@
Our specific contributions are listed below: For coded computing frameworks, we formulate a reconstruction methodology based on cubic B-spline interpolation. Our methodology explicitly specifies the construction of the spline space, the use of clamped knot vectors, and the natural boundary conditions, resulting in a well-posed linear system for reconstruction. We analyze the structure of the resulting B-spline interpolation system, and demonstrate that the reconstruction can be carried out efficiently using banded linear solvers, with computational complexity that scales linearly with the number of workers (see Section III). We establish explicit bounds on the operator norm associated with the proposed B-spline reconstruction and connect these bounds to the overall approximation error. This analysis yields a precise characterization of how the reconstruction error depends on the number of workers and stragglers. In particular, for a class of smooth non-polynomial functions, we show that the approximation error of BSCC exhibits an inverse-quadratic decay with respect to the number of workers (see Section IV). Finally, through experimental results, we show that BSCC outperforms BACC in terms of approximation accuracy for a broader class of smooth non-polynomial functions, however, with a marginal increase in the computational complexity during the reconstruction process (see Section V).
@@ II Preliminaries on B-Splines
In this section, we revisit some basic definitions and fundamental properties of B-spline functions used in the our construction and analysis. To define B-spline functions, we define knot vectors, which capture the support set of B-splines.
Definition 1.
A knot vector is a finite non-decreasing sequence of real numbers, denoted by $\mathbf{t}=\{t_{i}\}_{i=0}^{m-1}$, such that $t_{0}\leq t_{1}\leq\cdots\leq t_{m-1}$, where $m\in\mathbb{N}$. The elements of $\mathbf{t}$ are called knots.
Definition 2.
Let $p\geq 0$, and let $\mathbf{t}=\{t_{j}\}_{j=0}^{m-1}$ be a knot vector. Provided that $m\geq p+2$, the B-spline basis functions of degree $p$, denoted by $\{B_{j,p}(z)\mid j=0,1,\ldots,m-p-2\}$, are defined recursively over the knot vector $\mathbf{t}$ [11] as follows:
For $p=0$,
@@
If a denominator is zero, the corresponding term is defined to be zero.
Each B-spline basis function $B_{j,p}(z)$ satisfies $B_{j,p}(z)\geq 0$ for all $z$ and $B_{j,p}(z)=0$ for $z\notin[t_{j},t_{j+p+1})$. Consequently, each $B_{j,p}(z)$ has compact support and is nonzero only on the union of at most $p+1$ consecutive knot intervals. In the context of interpolation and approximation, B-spline basis functions are also defined over a special class of knot vectors called clamped knot vectors [12], which are formally defined below.
Definition 3.
Let $a,b]\subset\mathbb{R}$ be a compact interval and let $p\geq 0$ and $n>p$ be integers. Let $a=t_{0}<t_{1}<\cdots<t_{n-1}=b$ be a strictly increasing sequence of real numbers. A knot vector $\mathbf{t}$ is said to be $(p+1)$-clamped on $[a,b]$ if it is defined as
$\mathbf{t}=[\underbrace{t_{0},\ldots,t_{0}}_{p+1},t_{1},\ldots,t_{n-2},\underbrace{t_{n-1},\ldots,t_{n-1}}_{p+1}].$
Definition 4.
For $p\!\geq\!0$ and $m\!\geq\!p\!+\!2$, let $\{B_{j,p}(z)\mid j=0,1,\ldots,m\!-\!p\!-\!2\}$ denote the B-spline basis functions of degree $p$ defined with respect to a knot vector $\mathbf{t}=\{t_{i}\}_{i=0}^{m-1}$. Using these basis functions a spline space denoted by $\mathcal{S}_{p,\mathbf{t}}$, can be generated as[13]
$\mathcal{S}_{p,\mathbf{t}}=\bigg\{\sum_{j=0}^{m-p-2}C_{j,p}\,B_{j,p}(z)\;\forall\;C_{j,p}\in\mathbb{R}\bigg\}.$ (1)
The B-spline basis functions $\{B_{j,p}(z)\mid j=0,1,\ldots,m\!-\!p\!-\!2\}$ satisfy the partition of unity property, i.e., $\sum_{j=0}^{m-p-2}B_{j,p}(z)=1$, over its support set.
@@ III B-Spline Assisted Coded Computing
We propose a B-spline assisted coded computing (BSCC) framework for approximate computation of an arbitrary function $f(\cdot)$ over a given input dataset. Our setup consists of a master and $N$ workers indexed by the set $\mathcal{W}=\{\mathcal{W}_{0},\mathcal{W}_{1},\ldots,\mathcal{W}_{N-1}\}$. We assume that each worker communicates exclusively with the master through a dedicated communication link, and no communication is permitted among workers. The input dataset is stored at the master, while all workers are assumed to have prior knowledge of the function to be evaluated. Among the workers, we assume the presence of $S$ number of stragglers, for $1\leq S<N-2$, which may fail to return their computation results within the prescribed time. Let $\mathbf{X}=(\mathbf{X}_{0},\ldots,\mathbf{X}_{K-1})$ denote the dataset held by the master, where $\mathbf{X}_{j}\in\mathbb{R}^{m\times n}$ for all $j\in[K]$, with $K]\triangleq\{0,1,2,\ldots,K-1\}$. The primary objective of BSCC framework is to approximately evaluate an arbitrary function $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ on the matrices in $\mathbf{X}$ in a decentralized manner while guaranteeing tolerable approximation errors against $S$ number of stragglers. In particular, since the function $f(\cdot)$ may be non-polynomial, BSCC aims to enable numerically stable distributed computation of $f(\cdot)$ on $\mathbf{X}_{j}$ for all $j\in[K]$ with bounded approximation error; that is, if $\mathbf{Y}_{j}$ denotes the output reconstructed from the distributed computation corresponding to $\mathbf{X}_{j}$, then $\mathbf{Y}_{j}\approx f(\mathbf{X}_{j})$ for all $j\in[K]$ within a tolerable accuracy loss. Within this setting, we employ a generalized encoding method available in the literature and propose a B-spline assisted numerically stable reconstruction procedure at the master. [^1] Coded computing under additional constraints such as data privacy and Byzantine adversaries are well studied. In this work, we restrict our attention to resilience against stragglers. Privacy guarantees and Byzantine robustness are straightforward to generalize using existing techniques.
@@ III-A Encoding
We describe the encoding method, specifically highlighting how the dataset $\mathcal{X}$ is mapped into distributed shares for the workers. To encode the dataset, the master may employ any suitable interpolatory basis functions $\{\Phi_{j}(\cdot)\}$, $j\in[K]$, such as the Lagrange basis [4] or the Berrut basis functions [9]. These basis functions satisfy the cardinality property, namely $\Phi_{j}(\beta_{i})=1$ for $j=i$ and $\Phi_{j}(\beta_{i})=0$ for $j\neq i$, for the underlying $K$ ordered non-decreasing encoding points $\{\beta_{i}\in\mathbb{R}~|~i=0,1,\ldots,K-1\}$. Using this property, the dataset $\mathcal{X}$ is encoded as a function of an evaluation variable $z$ as
@@
III-B Distribution of Shares among the Workers
Once $u(z)$ is constructed as in (2), the master selects a set of $N$ distinct and ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}\subset\mathbb{R}$, such that $\alpha_{0}<\alpha_{1}<\cdots<\alpha_{N-1}$. For instance, these points could be chosen as Chebyshev nodes of the first kind or second kind. Using these evaluation points, the master generates the encoded shares by evaluating the encoding function $u(z)$ at each $\alpha_{i}$, resulting in the set of function values $\{\mathbf{Y}_{i}\}_{i=0}^{N-1}$, where $\mathbf{Y}_{i}=u(\alpha_{i}),\quad i\in[N].$ Finally, the master distributes the encoded shares to the workers by assigning the share $\mathbf{Y}_{i}$ to worker $\mathcal{W}_{i}$, for $i\in[N]$.
III-C Computation at the Workers
Once the worker $\mathcal{W}_{i}$ receives its evaluation $\mathbf{Y}_{i}$, it computes $f(\mathbf{Y}_{i})=f(u(\alpha_{i}))$, and returns it to the master. In contrast, the stragglers may not return these values within the prescribed deadline.
III-D Function Reconstruction using B-spline
In the absence of stragglers, i.e., when $S=0$, the master receives $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$ from $N$ workers, corresponding to the ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$. The objective of the master is to reconstruct an approximation of the composite function $f(u(z))$ over the interval spanned by these points. Since the function $f(\cdot)$ is arbitrary and may be non-polynomial, exact interpolation [6] is generally not feasible or numerically stable. Consequently, we approximate $f(u(z))$ using a B-spline function $r_{B-spline}(z)\approx f(u(z))$, where